How Many Litters of Carbon Dioxide gas is formed with 50 g of Calcium Carbonate decomposes?


1
Expert's answer
2021-02-15T03:09:05-0500

CaCO3 "\to" CaO + CO2

molar mass (M) of CaCO3 = 100.086 g/mol

no. of moles of CaCO3 = 50/100.086 = 0.499moles

CaCO3 and CaO and CO2 have equal moles

At STP of 273.15 K (0°C) and 1 atm, the molar volume of a gas is 22.414 L/mol.

Volume of CO2 = 0.499×22.414 = 11.18L